OpenStudy (anonymous):

If tan x = .88, what is cot x?

ganeshie8 (ganeshie8):

cot = 1/tan

ganeshie8 (ganeshie8):

simply flip the tan : \(\large \cot x = \dfrac{1}{\tan x} = \dfrac{1}{0.88} = ?\)